# Achieving Unbiased Multi-Instance Learning via Balanced Fine-Grained Positive-Unlabeled Learning

### v2

@@ 3. Derivation of Unbiased and Balanced PU Learning
Table 2. The table presents the AvgAcc of different algorithms under varying values of $\sigma_{\text{micro}}\in\{2,4,6,8,10\}$, along with the estimated AUC concerning changes in AvgAcc on the SST-2 Dataset.
@@
MIL ATT $80.76\pm 1.57$ $70.96\pm 3.72$ $57.49\pm 12.35$ $50.00\pm 0.00$ $51.16\pm 2.01$ $60.07$
FGSA $68.05\pm 16.87$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00\pm 0.00$ $53.61$
PReNET $78.82\pm 6.43$ $67.33\pm 3.25$ $66.49\pm 3.32$ $56.92\pm 6.07$ $60.47\pm 3.49$ $66.00$
@@
4. Balanced Fine-Grained PU Learning
4.1. Micro-to-Macro Optimization
(7) $\begin{split}F([y_{i1},\cdots,y_{il}])=\left\{\begin{array}[]{ll}+1,&\forall j\in\{1,\cdots,l\},y_{ij}=+1\\ -1,&\exists j\in\{1,\cdots,l\},y_{ij}=-1\end{array}\right.\end{split}$
(8) $G([g(x_{i1}),g(x_{i2}),\cdots,g(x_{il})])=\sum_{j=1}^{l}p(j=\arg\max_{k}g_{-1}(x_{ik}))g(x_{ij})$
Referring to eq. 8, we aim to find the idealized $p$ which makes $p(j=\arg\max_{k}g_{-1}(x_{ik}))=\mathbb{I}[j=\arg\max_{k}g_{-1}(x_{ik})]$. However, since we cannot obtain $g$, during the training process of neural networks, $\hat{g}$ from the neural network often deviates somewhat from the ideal decision function $g$. Consequently, we cannot directly obtain the value of the indicator function $\mathbb{I}[j=\arg\min_{k}g_{-1}(x_{ik})]$, but we can estimate the probability of the condition being satisfied. So, we use the probability function $\hat{p}$ instead of the function $p$.
(9) $\hat{p}(x_{ij})=\frac{\exp(\hat{g}_{-1}(x_{ij}))}{\sum_{k=1}^{l}\exp(\hat{g}_{-1}(x_{ik}))}$
$\hat{R}_{cgpn}(\hat{g})=$ $\frac{1}{2\cdot|P_{macro}|\cdot l}\sum_{X_{i}\in P_{macro}}\mathcal{L}[G([\hat{g}(x_{i1}),\cdots,\hat{g}(x_{il})]),+1]$
(10) $+$ $\frac{1}{2\cdot|N_{macro}|\cdot l}\sum_{X_{i}\in N_{macro}}\mathcal{L}[G([\hat{g}(x_{i1}),\cdots,\hat{g}(x_{il})]),-1]$
Based on eqs. 6, 8, 9 and 4.1, through the transformation of the learning paradigm, we derive a loss function that enables unbiased and balanced optimization of macro-level performance directly at the micro level which is called balanced fine-grained PU learning loss:
$\hat{R}_{cgpn}(\hat{g})=\hat{R}_{bfgpu}(\hat{g})=$ $\frac{1}{2|P_{micro}|}\sum_{x_{ij}\in P_{micro}}\hat{p}(x_{ij})\mathcal{L}[\hat{g}(x_{ij}),+1]$
(11) $+$ $\frac{1}{2|U_{micro}|}\sum_{x_{ij}\in U_{micro}}\hat{p}(x_{ij})\mathcal{L}[\hat{g}(x_{ij}),-1]$
4.2. Pseudo Labels Based on Macro Information
(12) $N_{pse}=\{(x_{ij},-1),j=\arg\max_{k}\hat{g}_{-1}(x_{ik}),X_{i}\in N_{macro}\}$
(13) $P_{pse}=\{(x_{ij},+1),j=\arg\min_{k}\hat{g}_{-1}(x_{ik}),X_{i}\in N_{macro}\}$
(14) $\hat{R}_{pse}(\hat{g})=\sum_{(x_{i},y_{i})\in N_{pse}+P_{pse}}\mathcal{L}[\hat{g}(x_{i}),y_{i}]$
4.3. Adjusted Decision Threshold (ADT)
(15) $T=sort([\hat{g}_{-1}(x_{i}),x_{i}\in U_{micro}])[\lfloor|U_{micro}|\cdot\pi\rfloor]$
This approach helps determine an accurate threshold because the decision function $g(x)$ reflects the conditional probability $p(y|x)$. $p(y|x)$ remains consistent between training and testing data. Therefore, the threshold obtained on the unlabeled data is highly applicable to the test data. The overall process of the algorithm can be seen in algorithm 1.
5. Experiments
5.1. Experimental Settings
All experiments were conducted using the Roberta model as the base model  (Liu et al., 2019), Adam as the optimizer  (Kingma and Ba, 2015), with a learning rate set to $10^{-5}$, and CosineAnnealingLR as the learning rate scheduler. The number of epochs was set to $5$, batch size to $16$, and all experiments were repeated $3$ times with random seeds $[0,1,2]$ to obtain the mean and variance of performance. We denote $\sigma_{micro}$ as the imbalance ratio at the micro level, which represents the positive-negative ratio of micro samples in negative macro samples. $\sigma_{macro}$ represents the imbalance ratio at the macro level, which is the ratio of positive macro samples to negative macro samples. We set $\lambda_{bfgpu}=1/\pi$ and $\lambda_{pse}=1$ in all experiments. For public datasets, a macro-level normal sample is composed of $\sigma_{micro}+1$ positive instances, while a macro-level anomalous sample is composed of $\sigma_{micro}$ positive instances and $1$ negative instanceIn the experiments in sections 5.3, 5.4 and 5.6, $\sigma_{macro}$ was set to $1$, while in the experiments in section 5.5, $\sigma_{macro}$ was set to $5$. $\sigma_{micro}$ was set to $[2,4,6,8,10]$ for short texts and $[2,3,4,5]$ for moderate texts. In section 5.4, $\sigma_{micro}$ varied inconsistently averaging around $25$. For micro-level PU learning, $\pi=\frac{\sigma_{micro}}{\sigma_{micro}+1}$. The algorithms were implemented using the PyTorch framework  (Paszke et al., 2019). We used average accuracy and F1 score, two commonly used evaluation metrics in imbalanced learning. Regarding the F1 score, when dealing with class imbalance problems, we typically focus on the performance of the minority (anomalous) class. We also compared the algorithm performance as $\sigma_{micro}$ varies, using the area under the curve $AUC_{AvgAcc}$ and $AUC_{F1}$. 4 A800 GPUs are used for all experiments.
5.2. Compared Methods
Our comparative algorithms consist of: Macro PN: Conventional macro binary classification. Macro DeepSAD: Supervised macro AD using DeepSAD  (Ruff et al., 2019). Micro DeepSAD: Supervised micro AD, using DeepSAD  (Ruff et al., 2019) and treating unlabeled samples as negative class. Micro DeepSVDD: Unsupervised micro AD, using DeepSVDD  (Ruff et al., 2018) which is an One-Class Classification (OCC) method using only positive samples. Micro Multi-Instance Learning: Multi-instance learning algorithms used to address inexact supervision, using three multi-instance learning algorithms MIL-ATT  (Ilse et al., 2018), MIL-FGSA  (Angelidis and Lapata, 2018), and MIL-PReNET (Pang et al., 2023) which are suitable for text data. Micro PU Learning: Utilizing four types of loss functions uPU  (Du Plessis et al., 2015), nnPU  (Kiryo et al., 2017), balancedPU  (Su et al., 2021), and robustPU (Zhu et al., 2023) for micro-level PU learning.
Figure 3. The figure presents the AvgAcc of different algorithms under both micro and macro imbalance on SST-2.
Table 3. Experiments on Customer Service Quality Inspection Dataset in the Real Application.
Method AvgAcc F1 Score
Macro PN $52.38\pm 4.12$ $8.33\pm 14.33$
DeepSAD $47.61\pm 4.12$ $6.67\pm 11.55$
MIL ATT $54.76\pm 11.48$ $53.27\pm 8.72$
FGSA $53.57\pm 7.14$ $54.14\pm 9.12$
PReNET $59.52\pm 2.06$ $59.05\pm 0.86$
Micro DeepSAD $58.33\pm 5.46$ $63.12\pm 4.89$
DeepSVDD $50.00\pm 0.00$ $00.00\pm 0$ 0.00
uPU $50.00\pm 0.00$ $00.00\pm 0$ 0.00
nnPU $50.00\pm 0.00$ $00.00\pm 0$ 0.00
balancedPU $50.00\pm 0.00$ $00.00\pm 0$ 0.00
robustPU $50.00\pm 0.00$ $00.00\pm 0$ 0.00
BFGPU $71.43\pm 9.45$ $75.31\pm 5.84$
5.3. Experiments with Public Datasets
We conducted experiments with 2 datasets with small imbalanced ratios: IMDB (Maas et al., 2011), Amazon, and 2 with large ones: SST-2  (Socher et al., 2013), Sentiment140  (Go et al., 2009). We present the AvgAcc results on the SST-2 and IMDB datasets in tables 2 and 1. The remaining experimental results can be found in tables 5, 4, 7 and 6, which are in the appendix.
5.4. Experiments in the Real Application
In a real-world application, we evaluated with a customer service quality inspection dataset depicted in fig. 1. The dataset consists of instances where service personnel were flagged for substandard performance during service. The results are shown in table 3.
5.5. Experiments with Imbalance at Both Level
@@
We considered a more extreme scenario where the normal-anomalous ratio at not only the micro but also the macro level is imbalanced. The imbalance lead to the anomalous information accounting for only $\frac{1}{(1+\sigma_{micro})\cdot(1+\sigma_{macro})}$. However, our algorithm still achieved outstanding performance. We set $\sigma_{macro}$ to 5 and conducted the experiment on the SST-2 dataset. The experimental results are shown in table 8 in the appendix. We plotted the curve of average accuracy varying as shown in fig. 3.
5.6. Ablation Study
@@
To validate the necessity of each component of BFGPU, we conducted ablation experiments. In addition to the comprehensive algorithm, we considered 3 settings: Not using the balanced fine-grained PU learning loss function; Not using pseudo-labeling for further training; Not using adjusted thresholds. The results on the datasets IMDB and SST-2 are shown in table 9 and table 10 respectively.
5.7. Sensitivity Analysis
@@
BFGPU mainly has two hyperparameters, $\lambda_{bfgpu}$ and $\lambda_{pse}$. To validate the stability under different hyperparameter settings, we conducted a sensitivity analysis using the SST-2 dataset. The results of $\lambda_{bfgpu}$ and $\lambda_{pse}$ are shown in table 11 and table 12 respectively.
5.8. Analysis of Results
All results pointed to some common conclusions: balanced fine-grained PU learning outperformed coarse-grained PN learning and MIL; in most cases, BFGPU is more suitable for the MIL problem than other PU methods; in real-world and extremely imbalanced scenarios, BFGPU is irreplaceable. The ablation experiments demonstrate that the absence of each component of the algorithm framework significantly reduces the overall performance. The sensitivity analysis shows that BFGPU performs sufficiently well and is less dependent on the hyperparameters.
6. Theoretical Argumentation
@@
(17) $+\frac{\sigma_{macro}+1}{2}\cdot\sqrt{\frac{2\cdot(\sigma_{micro}+1)\cdot ln(4/\delta)}{|U_{micro}|}}+Disc(\mathcal{G},p_{1}(x,y),p_{1/l}(x,y))$
@@
(18) $Disc(\mathcal{G},p_{1}(x,y),p_{1/l})=\max_{g\in\mathcal{G}}|p_{x,y\sim\ p_{1}(x,y)}(g(x)\not=y)-p_{x,y\sim\ p_{1/l}(x,y)}(g(x)\not=y)|$
@@
(19) $\cdot$ $\sqrt{\frac{2\cdot ln(4/\delta)}{|P_{micro}|}}+\frac{\sigma_{macro}+1}{2}\cdot\sqrt{\frac{2\cdot ln(4/\delta)}{|U_{micro}|}}+p_{inc}(\mathcal{G},g^{*})+\frac{\sigma_{macro}+1}{2}\cdot\frac{\sigma_{micro}}{\sigma_{micro}+1}$
@@
Based on the derivation of the PU loss in section 4.1, we found that directly optimizing unlabeled samples as abnormal samples essentially balances the loss of PN learning. Therefore, the generalization error bound of BFGPU that we derive is:
@@
(20) $\hat{R}(g_{bfgpu})-R(g^{*})$ $\leq\frac{4\cdot\alpha_{L}\cdot C_{\mathcal{G}}}{\sqrt{|P_{micro}|}}+\sqrt{\frac{2\cdot ln(4/\delta)}{|P_{micro}|}}+\frac{4\cdot\alpha_{L}\cdot C_{\mathcal{G}}}{\sqrt{|U_{micro}|}}+\sqrt{\frac{2\cdot ln(4/\delta)}{|U_{micro}|}}+p_{inc}(\mathcal{G},g^{*}).$
@@
7. Conclusion
We addressed the challenge of anomalous detection in real-world scenarios where anomalous information is extremely sparse: macroscopically, there is high similarity between normal and anomalous samples, and microscopically, labels are lacking and imbalanced. We transformed the problem into an imbalanced PU learning task at the micro level, demonstrating its feasibility through theoretical analysis. We proposed a solution that directly optimizes macro-level performance at the micro level using PU learning objectives and combined pseudo-labeling with threshold adjustment techniques to create a new framework. Experiments validated that this framework effectively tackles the issue of sparse negative information while maintaining strong performance even in extreme cases of imbalance at both macro and micro levels.
Appendix A Algorithm Framework
The specific algorithmic framework is shown in algorithm 1.
@@ Appendix A Algorithm Framework
        Get the probabilities $\hat{p}$ by eq. 9
        Get the loss $\hat{R}_{bfgpu}$ by section 4.1
@@
Appendix B Proof of the Theories
B.1. Proof of Theorem 4.1
@@ Appendix B Proof of the Theories
(21) $+$ $\frac{1}{2}\cdot(\frac{4\cdot\alpha_{L}\cdot C_{\mathcal{G}}}{\sqrt{|U_{micro}|/(\sigma_{micro}+1)}}+\sqrt{\frac{2\cdot ln(4/\delta)}{|U_{micro}|/(\sigma_{micro}+1)}})/\frac{1}{\sigma_{macro}+1}$
@@
(23) $+$ $\frac{\sigma_{macro}+1}{2}\cdot\sqrt{\frac{2\cdot(\sigma_{micro}+1)\cdot ln(4/\delta)}{|U_{micro}|}}+Disc(\mathcal{G},p_{1}(x,y),p_{1/l}(x,y))$
B.2. Proof of Theorem 4.2
@@
(24) $+$ $\frac{1}{2}\cdot(\frac{4\cdot L_{l}\cdot C_{\mathcal{G}}}{\sqrt{|U_{micro}|}}+\sqrt{\frac{2\cdot ln(4/\delta)}{|U_{micro}|}}+\frac{\sigma_{micro}}{\sigma_{micro}+1})/\frac{1}{\sigma_{macro}+1}$
@@
(26) $+$ $\frac{\sigma_{macro}+1}{2}\cdot\sqrt{\frac{2\cdot ln(4/\delta)}{|U_{micro}|}}+p_{inc}(\mathcal{G},g^{*})+\frac{\sigma_{macro}+1}{2}\cdot\frac{\sigma_{micro}}{\sigma_{micro}+1}$
B.3. Proof of Theorem 4.3
@@
(27) $+$ $\sqrt{\frac{2\cdot ln(4/\delta)}{|P_{micro}|}}+\sqrt{\frac{2\cdot ln(4/\delta)}{|U_{micro}|}}$
@@
(29) $+$ $\sqrt{\frac{2\cdot ln(4/\delta)}{|P_{micro}|}}+\sqrt{\frac{2\cdot ln(4/\delta)}{|U_{micro}|}}+p_{inc}(\mathcal{G},g^{*})$
Appendix C Additional Experimental Results
C.1. Experiments with Low Imbalance Ratio
We present a comparative experiment on two long text datasets: IMDB and Amazon. The parameter $\sigma_{micro}$ is set to [2, 3, 4, 5]. Due to space limitations in the main text, the additional experiments on the datasets Amazon are shown in table 6 in this appendix.
Table 4. The table presents the F1 scores of different algorithms under varying values of $\sigma_{\text{micro}}\in\{2,3,4,5\}$, along with the estimated AUC with respect to changes in F1 on the IMDB Dataset.

### v3

@@ 5.5 Experiments with Synthetic Datasets
Table 3: The table presents the AvgAcc of various algorithms under varying values of $\sigma_{\text{micro}}\in\{2,4,6,8,10\}$, along with the estimated AUC concerning changes in AvgAcc on the SST-2 Dataset.
@@
MIL MaxPooling $85.82\pm 1.78$ $66.83\pm 6.44$ $61.84\pm 10.31$ $52.83\pm 4.90$ $56.20\pm 10.74$ $64.70$
TopkPooling - $75.25\pm 5.35$ $75.85\pm 2.74$ $68.87\pm 3.27$ $64.73\pm 12.81$ $71.18$
Attention $80.76\pm 1.57$ $70.96\pm 3.72$ $57.49\pm 12.35$ $50.00\pm 0.00$ $51.16\pm 2.01$ $60.07$
FGSA $68.05\pm 16.87$ $\pm 0.00 50.00±0.005$$\pm 0.00$ $\pm 0.00 50.00±0.005$$\pm 0.00$ $53.61$
PReNET $78.82\pm 6.43$ $7.33\pm 3.25 66$$6.49\pm 3.32$ $56.92\pm 6.07$ $47\pm 3.49 66.00$66.00
@@
We considered a more extreme scenario where the normal-anomalous ratio at not only the micro but also the macro level is imbalanced. This lead to the anomalous information accounting for only $\frac{1}{(1+\sigma_{micro})\cdot(1+\sigma_{macro})}$. However, BFGPU still achieved outstanding performance. We set $\sigma_{macro}$ to 5 and 10 and conducted the experiments. We plotted the curve of average accuracy varying using synthetic SST-2 as shown in fig 3 and 4. The experimental results are shown in tables˜9 and 10.
Figure 3: The figure presents the AvgAcc of various algorithms when $\sigma_{macro}=5$.
Figure 4: The figure presents the AvgAcc of various algorithms when $\sigma_{macro}=10$.
5.6 Compared with LLMs
We maximized the $\sigma_{micro}$ in the synthetic datasets, setting it to 5 for IMDB and Amazon, and to 10 for SST-2 and Sentiment140, to evaluate whether the model can accurately identify samples with slight anomalies. This further enabled a comparison of the "needle-in-a-haystack" capability between BFGPU and current mainstream LLMs. We used a consistent prompt: "You are a sentiment classification model; output ‘positive’ if there is no negative sentiment in the paragraph, and ‘negative’ if negative sentiment exists. You must only output ‘positive’ or ‘negative’ without any additional content." The experimental results, shown in table˜4, confirm that BFGPU is currently more suitable than LLMs themselves as a reward model in the RLHF process.
Table 4: The table presents the performance compared with LLMs with maximum $\sigma_{micro}$.
Model IMDB Amazon SST-2 Sentiment140
AvgAcc F1 Score AvgAcc F1 Score AvgAcc F1 Score AvgAcc F1 Score
GPT-4-turbo 64.71 71.43 79.41 86.27 70.45 74.95 55.88 59.46
Qwen3 67.65 66.67 82.35 86.36 67.86 64.52 64.71 62.50
Deepseek v3 55.88 48.28 67.65 71.79 57.58 56.25 61.76 55.17
Gemini2.5 64.71 62.50 79.41 85.11 63.09 74.42 58.82 61.11
BFGPU $83.64$ $84.61$ $85.57$ $86.50$ $82.56$ $84.41$ $66.45$ $66.40$
5.7 Ablation Study and Sensitivity Analysis
@@
We conducted some ablation experiments. a) To validate the necessity of each component of BFGPU, we considered 3 settings: Not using the BFGPU learning loss function; Not using pseudo-labeling for further training; Not using adjusted thresholds. The results are shown in tables˜11 and 12. b) To verify that our results are not limited by the choice of backbone model, we conducted further evaluations using DeBERTa table˜13. The results are shown in table˜13. c) To further demonstrate the superiority of the BFGPU loss function, we compared its performance against other PU learning methods when integrated with our proposed pseudo-labeling and ADT strategies. The results are shown in table˜14.
@@
BFGPU mainly has two hyperparameters, $\lambda_{bfgpu}$ and $\lambda_{pse}$. To validate the stability under different hyperparameter settings, we conducted a sensitivity analysis using the SST-2 dataset. The results of $\lambda_{bfgpu}$ and $\lambda_{pse}$ are shown in table˜15 and table˜16 respectively.
6 Theoretical Argumentation
@@
$+\frac{\sigma_{macro}+1}{2}\cdot\sqrt{\frac{2\cdot(\sigma_{micro}+1)\cdot ln(4/\delta)}{|U_{micro}|}}+Disc(\mathcal{G},p_{1}(x,y),p_{1/l}(x,y))$ (17)
@@
$Disc(\mathcal{G},p_{1}(x,y),p_{1/l})=\max_{g\in\mathcal{G}}|p_{x,y\sim\ p_{1}(x,y)}(g(x)\not=y)-p_{x,y\sim\ p_{1/l}(x,y)}(g(x)\not=y)|$ (18)
@@
$\cdot$ $\sqrt{\frac{2\cdot ln(4/\delta)}{|P_{micro}|}}+\frac{\sigma_{macro}+1}{2}\cdot\sqrt{\frac{2\cdot ln(4/\delta)}{|U_{micro}|}}+p_{inc}(\mathcal{G},g^{*})+\frac{\sigma_{macro}+1}{2}\cdot\frac{\sigma_{micro}}{\sigma_{micro}+1}$ (19)
@@
Based on the derivation of the PU loss in section˜4.1, we found that directly optimizing unlabeled samples as abnormal samples essentially balances the loss of PN learning. Therefore, the generalization error bound of BFGPU that we derive is:
@@
$\hat{R}(g_{bfgpu})-R(g^{*})$ $\leq\frac{4\cdot\alpha_{L}\cdot C_{\mathcal{G}}}{\sqrt{|P_{micro}|}}+\sqrt{\frac{2\cdot ln(4/\delta)}{|P_{micro}|}}+\frac{4\cdot\alpha_{L}\cdot C_{\mathcal{G}}}{\sqrt{|U_{micro}|}}+\sqrt{\frac{2\cdot ln(4/\delta)}{|U_{micro}|}}+p_{inc}(\mathcal{G},g^{*})$ (20)
@@
7 Conclusion
We addressed the challenge of scarce and sparse anomalous detection in real-world scenarios: macroscopically, there is high similarity between normal and anomalous samples, and microscopically, labels are lacking and imbalanced. Accordingly, we propose a dual-imbalanced MIL problem. We transformed the problem into an imbalanced PU learning task at the micro level, demonstrating its feasibility through theoretical analysis. We proposed a solution that directly optimizes macro-level performance at the micro level using PU learning objectives and combined pseudo-labeling with threshold adjustment techniques to create a new framework. Experiments validated that this framework effectively tackles the issue of sparse negative information while maintaining strong performance even in extreme cases of imbalance at both macro and micro levels.
Ethics Statement
We foresee no direct negative societal impacts resulting from this work, which is intended to advance research in computational music and creativity.
Reproducibility Statement
We are fully committed to the reproducibility of our research. The code for this work has been open-sourced at https://github.com/BFGPU/BFGPU, and the experimental setup section of the paper provides a comprehensive description of all models and parameters used. We will subsequently focus on providing a more user-friendly interface to facilitate their use of the proposed method as much as possible.
Appendix A LLMs Usage Statement
We used Large Language Models (LLMs) to assist with the writing of this paper. Their primary role was to improve grammar, phrasing, and clarity.
Appendix B Algorithm Framework
This section presents the specific algorithmic framework.
@@ Appendix B Algorithm Framework
   Get the probabilities $\hat{p}$ by eq.˜9
   Get the loss $\hat{R}_{bfgpu}$ by section˜4.1
@@
Appendix C Proof of the Theories
C.1 Proof of Theorem 4.1
@@ Appendix C Proof of the Theories
$+$ $\frac{1}{2}\cdot(\frac{4\cdot\alpha_{L}\cdot C_{\mathcal{G}}}{\sqrt{|U_{micro}|/(\sigma_{micro}+1)}}+\sqrt{\frac{2\cdot ln(4/\delta)}{|U_{micro}|/(\sigma_{micro}+1)}})/\frac{1}{\sigma_{macro}+1}$ (21)
@@
$+$ $\frac{\sigma_{macro}+1}{2}\cdot\sqrt{\frac{2\cdot(\sigma_{micro}+1)\cdot ln(4/\delta)}{|U_{micro}|}}+Disc(\mathcal{G},p_{1}(x,y),p_{1/l}(x,y))$ (23)
C.2 Proof of Theorem 4.2
@@
$+$ $\frac{1}{2}\cdot(\frac{4\cdot L_{l}\cdot C_{\mathcal{G}}}{\sqrt{|U_{micro}|}}+\sqrt{\frac{2\cdot ln(4/\delta)}{|U_{micro}|}}+\frac{\sigma_{micro}}{\sigma_{micro}+1})/\frac{1}{\sigma_{macro}+1}$ (24)
@@
$+$ $\frac{\sigma_{macro}+1}{2}\cdot\sqrt{\frac{2\cdot ln(4/\delta)}{|U_{micro}|}}+p_{inc}(\mathcal{G},g^{*})+\frac{\sigma_{macro}+1}{2}\cdot\frac{\sigma_{micro}}{\sigma_{micro}+1}$ (26)
C.3 Proof of Theorem 4.3
@@
$+$ $\sqrt{\frac{2\cdot ln(4/\delta)}{|P_{micro}|}}+\sqrt{\frac{2\cdot ln(4/\delta)}{|U_{micro}|}}$ (27)
@@
$+$ $\sqrt{\frac{2\cdot ln(4/\delta)}{|P_{micro}|}}+\sqrt{\frac{2\cdot ln(4/\delta)}{|U_{micro}|}}+p_{inc}(\mathcal{G},g^{*})$ (29)
Appendix D Additional Experimental Results
D.1 Experiments with Synthetic Datasets
Due to space limitations in the main text, the additional experiments on the SST-2 and Sentiment140 datasets are provided in tables˜5 and 6 in this appendix.
Table 5: The table presents the F1 scores of various algorithms under varying values of $\sigma_{\text{micro}}\in\{2,4,6,8,10\}$, along with the estimated AUC with respect to changes in F1 on the SST-2 Dataset.
Methods F1 Score $AUC_{F1}$
2 4 6 8 10
Macro Supervised $83.63\pm 0.21$ $79.35\pm 1.04$ $78.15\pm 2.81$ $75.77\pm 5.79$ $75.50\pm 3.23$ $78.48$
MIL MaxPooling $85.28\pm 7.74$ $52.34\pm 4.00$ $37.42\pm 3.35$ $12.38\pm 4.36$ $19.49\pm 4.03$ $41.38$
TopkPooling - $69.95\pm 6.85$ $70.39\pm 4.89$ $61.75\pm 4.28$ $44.44\pm 38.49$ $61.63$
Attention $80.54\pm 0.51$ $62.39\pm 7.22$ $45.00\pm 36.50$ $0.00\pm 0.00$ $48.79\pm 30.96$ $47.34$
FGSA $9.38\pm 43.89 44$$4.44\pm 38.49$ $66.67\pm 0.00$ $44\pm 38.49 44.44$$44\pm 38.49$ $49.87$
Micro DeepSAD $51.95\pm 12.30$ $14.95\pm 2.18$ $01\pm 7.18 9.9$$92\pm 2.40$ $14.18\pm 4.29$ $20.00$
uPU $76.82\pm 3.40$ $64.43\pm 6.62$ $62.13\pm 6.88$ $60.79\pm 10.15$ $7.56\pm 1.27 66$.3566.35
nnPU $84.00\pm 1.25$ $6.74\pm 0.10 66$$67\pm 0.00 66.$$67\pm 0.00 66.$$6.67\pm 0.00$ $70.15$
Table 6: The table presents AvgAcc and F1 scores of various algorithms under varying values of $\sigma_{\text{micro}}\in\{2,4,6,8,10\}$, along with the estimated AUC with respect to changes in AvgAcc and F1 on the Sentiment140 Dataset.
Method AvgAcc $AUC_{AvgAcc}$
2 4 6 8 10
Macro Supervised $74.37\pm 0.40$ $9.80\pm 0.37 66$$6.81\pm 0.48$ $65.09\pm 0.19$ $62.20\pm 0.48$ $67.65$
MIL MaxPooling $60.79\pm 2.63$ $62.51\pm 9.54$ $56.78\pm 5.84$ $49.95\pm 0.09$ $52.03\pm 3.41$ $56.41$
TopkPooling - $65.08\pm 1.90$ $62.14\pm 2.28$ $56.31\pm 5.96$ $52.95\pm 5.07$ $59.12$
Attention $71.27\pm 0.36$ $64.97\pm 3.24$ $52.88\pm 4.98$ $50.06\pm 0.10$ $50.16\pm 0.28$ $57.87$
FGSA $64.17\pm 3.13$ $53.46\pm 4.03$ $\pm 0.00 50.00±0.005$$\pm 0.00$ $50.00\pm 0.00$ $53.53$
PReNET $70.41\pm 3.74$ $7.11\pm 7.11 55$$5.56\pm 4.83$ $54.85\pm 2.93$ $50.00\pm 0.00$ $57.59$
Micro DeepSAD $51.07\pm 2.14$ $51.80\pm 2.05$ $65.23\pm 2.04$ $63.99\pm 8.28$ $\pm 8.64 58.1858.18$
nnPU $69.66\pm 1.01$ $\pm 0.00 50.00±0.005$$\pm 0.00$ $\pm 0.00 50.00±0.005$$\pm 0.00$ $53.93$
robustPU $74.65\pm 0.70$ $70.13\pm 1.63$ $66.94\pm 0.40$ $59.14\pm 8.58$ $2.88\pm 0.27 66$.7466.74
BFGPU $75.86\pm 1.01$ $70.68\pm 0.97$ $7.95\pm 1.53 66$$6.34\pm 1.25$ $64.45\pm 0.35$ $69.06$
Method F1 Score $AUC_{F1}$
2 4 6 8 10
Macro Supervised $73.99\pm 0.72$ $68.75\pm 0.44$ $64.94\pm 1.66$ $63.86\pm 0.52$ $59.50\pm 1.48$ $66.21$
MIL MaxPooling $37.85\pm 7.23$ $43.23\pm 30.95$ $27.71\pm 23.78$ $21.92\pm 37.96$ $9.96\pm 16.53$ $28.13$
TopkPooling - $52.83\pm 4.60$ $47.32\pm 4.56$ $26.12\pm 23.20$ $34.85\pm 33.37$ $40.28$
Attention $67.64\pm 0.52$ $56.56\pm 5.58$ $37.40\pm 34.07$ $22.85\pm 37.96$ $23.34\pm 37.56$ $41.56$
FGSA $52.19\pm 8.97$ $16.28\pm 17.59$ $44\pm 38.49 44.44$$\pm 38.49 44.44$$44\pm 38.49$ $40.36$
Micro DeepSAD $4.65\pm 1.80 55$$5.89\pm 1.76$ $7.50\pm 1.80 66$$6.55\pm 7.42$ $61.86\pm 7.61$ $61.29$
uPU $60.88\pm 2.42$ $57.32\pm 10.29$ $48.65\pm 13.84$ $0.66\pm 4.54 66$$6.66\pm 0.21$ $58.83$
nnPU $75.50\pm 0.56$ $6.67\pm 0.00 66$$67\pm 0.00 66.$$67\pm 0.00 66.$$6.67\pm 0.00$ $68.44$
BFGPU $77.21\pm 0.73$ $73.27\pm 0.95$ $69.80\pm 3.18$ $7.21\pm 2.94 66$$6.40\pm 0.68$ $\bm{70.78}$
We present a comparative experiment on two long text datasets: IMDB and Amazon. The parameter $\sigma_{micro}$ is set to [2, 3, 4, 5]. Due to space limitations in the main text, the additional experiments on the datasets Amazon are shown in tables˜7 and 8 in this appendix.
Table 7: The table presents the AvgAcc and F1 scores of various algorithms under varying values of $\sigma_{\text{micro}}\in\{2,3,4,5\}$, along with the estimated AUC on the IMDB Dataset.
Method AvgAcc $AUC_{AvgAcc}$
2 3 4 5
Macro Supervised $76.75\pm 1.45$ $70.62\pm 0.83$ $62.61\pm 4.70$ $58.33\pm 4.81$ $67.08$
DeepSAD $52.17\pm 0.56$ $2.57\pm 1.20 55$$5.01\pm 1.05$ $50.76\pm 0.77$ $52.63$
MIL MaxPooling $85.76\pm 2.55$ $74.30\pm 18.50$ $60.31\pm 15.71$ $55.56\pm 9.63$ $68.98$
TopkPooling - $82.23\pm 1.22$ $83.27\pm 0.61$ $82.41\pm 2.17$ $82.63$
Attention $82.58\pm 7.37$ $85.98\pm 1.32$ $80.66\pm 3.81$ $78.62\pm 1.54$ $81.96$
PReNET $84.53\pm 1.50$ $80.54\pm 1.50$ $71.78\pm 4.97$ $3.56\pm 2.49 77$.6077.60
uPU $82.93\pm 1.14$ $76.30\pm 1.51$ $78.57\pm 5.37$ $2.29\pm 3.30 77$.5277.52
nnPU $84.53\pm 1.49$ $58.39\pm 5.95$ $\pm 0.00 50.00±0.005$$\pm 0.00$ $60.73$
BFGPU $88.13\pm 1.07$ $87.93\pm 0.58$ $86.02\pm 0.69$ $\pm 0.78 86.4386.43$